OpenStudy (anonymous):

Logarithms Solve 6^2x-2(6^x)-15=0

OpenStudy (anonymous):

Say m=6^x M^2-2m-15=0 (m-5)(m+3)=0 m=5 or -3 So 6^x= 5 or -3 C

OpenStudy (anonymous):

\[x= {\log_{6} 5}\] or\[{ \log_{6}-3}\]

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

Beware..the second answer is extraneous. You can't take a log of a negative value. (Think about it: 6 to what power becomes suddenly negative?)

OpenStudy (anonymous):

Ahh yeah thanks